20. The hourly wage for an employee is 11$ per hour plus an additional $.80 for every unit produced during the hour. The employee earned a total of $23 during the first hour.How many units were produced in that hour?
Can someone help with this equation?

Answer by edjones(8007)

Let x=number of units produced during an hour
11+.8x=23
subtract 11 from both sides. .8x=12
divide .8 into both sides. x=15 units
Check:
11+(.8*15)=11+12=23
Ed